OpenStudy (anonymous):

how to solve xy''+2y'+xy=0

OpenStudy (anonymous):

you sure this is the given?

OpenStudy (nikvist):

\[xy''+2y'+xy=0\]\[(xy)'=y+xy'\]\[(xy)''=(y+xy')'=y'+y'+xy''=2y'+xy''\] \[(xy)''+xy=0\quad,\quad z=xy\quad,\quad z''+z=0\]\[z=C_1\cos{x}+C_2\sin{x}\quad\Rightarrow\quad y=\frac{z}{x}=\frac{C_1\cos{x}+C_2\sin{x}}{x}\]